OpenStudy (anonymous):

Solve the inequality. -5.3 > 6.7 + 4.3 + q -

OpenStudy (anonymous):

There are two cases, when it is \[\ge\] First case: \[-5.3\ge6.7+4.3+q\] Adding 6.7 and 4.3 gives us 11, take 11 to other side it becomes -11, so: \[-5.3-11\ge q\] \[-16.3\ge q\] That's the first case. The second case is: \[-5.3\le-6.7-4.3-q\] \[-5.3\le-11-q\] \[-5.3+11\le-q\] \[5.7\le -q\]